OpenStudy (anonymous):

prove that for every n>=1, the number 11^n+1 + 12^2n-1 is divisible by 133. ( using math induction)

OpenStudy (anonymous):

Prove for n=1 first, to see whether it is true.

OpenStudy (anonymous):

Assume the formula is true for n=k. Write that expression in terms of k. Then prove for n=k+1.

OpenStudy (anonymous):

Which step are you stuck at the moment?

OpenStudy (anonymous):

i know how to do the first two step. its just the last part

OpenStudy (anonymous):

it just takes factoring

OpenStudy (anonymous):

Note: \(133+11=144\)

OpenStudy (anonymous):

hence \(144-11=12^2-11=133\)... do you see what to do now?

OpenStudy (anonymous):

I tried doing thid: 133/ 11^k+2 + 12^2k+1 = 11*11^k+1 +12^2 * 12^2k-1

OpenStudy (anonymous):

Presume \(11^{n+1}+12^{2n-1}=133k\) for some integers \(n,k\).

OpenStudy (anonymous):

Now observe:$$11^{(n+1)+1}+12^{2(n+1)-1}=11\times11^{n+1}+12^2\times12^{2n-1}$$

OpenStudy (anonymous):

$$11\times11^{n+1}+144\times12^{2n-1}=11\times11^{n+1}+(11+133)\times12^{2n-1}=\cdots$$

OpenStudy (anonymous):

Using what oldrin.bataku wrote down above, you prove: For n=k+1. \[\large LHS=11^{k+2}+12^{2k+1}\] \[\large =11(11^{k+1})+12^{2}(12^{2k+1})\] Then you factor out the 11 and ignore the second term and copy down what you wrote for n=k.

OpenStudy (anonymous):

By doing that, you have to RIG the expression.

OpenStudy (anonymous):

$$11\times11^{n+1}+11\times 12^{2n-1}+133\times12^{2n-1}=11\times(11^{n+1}+12^{2n-1})+133\times12^{2n-1}$$

OpenStudy (anonymous):

so$$11\times133k+133\times12^{2n-1}=133(11k+12^{2n-1})$$ which is clearly divisible by \(13\)

OpenStudy (anonymous):

133*

OpenStudy (anonymous):

Couldn't have said it better myself @oldrin.bataku

OpenStudy (anonymous):

so i first do the base cake n=1 133/ 11^2 + 12 133/ 133 true assume n = k 133/ 11^k+1 +12^2k-1 tru

OpenStudy (anonymous):

Why are you assuming n = k... just presume there exists some \(n\) such that our conjecture holds and then show \(n+1\) also holds.

OpenStudy (anonymous):

No you should write an expression. Write a formula like oldrin did then.

OpenStudy (anonymous):

shouldn't*

OpenStudy (anonymous):

I thought that one has to first do the base case, then assume true for n=k and then prove for n= k+1

OpenStudy (anonymous):

And it's better to assume that n=k oldrin. SO you don't get confused with n=n.

OpenStudy (anonymous):

@mathtutoring22 That's a good way of doing mathematical induction. Continue doing that.

OpenStudy (anonymous):

@mathtutoring22 the \(k\) I used in mine is indeterminate... it's just there to state that our integer is divisible by \(133\)

OpenStudy (anonymous):

When you assume the formula is true for n=k, you should write this down: \[11^{k+1} +12^{2k-1}=133m\] where m is an integer.

OpenStudy (anonymous):

oh okay that's fine

OpenStudy (anonymous):

You can use any old letter. Just state that it is an integer etc. when you do.

OpenStudy (anonymous):

how should i put this all together?

OpenStudy (anonymous):

Well you should put all that chronologically and then show us so we can instantly tell what you need to fix etc.

OpenStudy (anonymous):

@mathtutoring22 Presume for some positive integer \(k\) we have $$11^{k+1}+12^{2k-1}=133m$$for some integer \(m.\) Observe:$$11^{(k+1)+1}+12^{2(k+1)-1}=11\times11^{k+1}+12^2\times12^{2k-1}=\dots$$Clearly, \(11^{1+1}+12^{2(1)-1}=11^2+12=121+12=133\) is divisible by \(133\). Hence \(11^{k+1}+12^{2k-1}\) is divisible by \(133\) for all \(k\ge1\). as an example...

OpenStudy (anonymous):

Base Case?

OpenStudy (anonymous):

idk why im so lost

OpenStudy (anonymous):

I don't like following a rigid proof template... just make sure the logic checks out

OpenStudy (anonymous):

@mathtutoring22 Just write down everything for this proof and we will tell you what you did wrong and we can patch things up for you. WE can't do that unless you show us what you've done.

OpenStudy (anonymous):

like Base case: .... Assume true for n= k ... Prove true for n= k+1

OpenStudy (anonymous):

And then write your conclusion. You've got everything in the right order. I don't see anything wrong with that. So why are you making a fuss of being confused?

OpenStudy (anonymous):

this is what i did: Base case: 133/ 11^2 +12 133/133 true Assume true for n=k 133/ 11^(k+1) +12^2k-1 prove for n= K+1 133/ 11^k+2 + 12^2k+2 = 11*11^k+1 + 12^2 * 12^2k-1 = 11*11^11^k+1 + (133+11)*12^12^2k-1 = 11*11^k+1*133*^12^2k-1 + 11*12^2k-1 = 11*(11^K+1+12^2k-1) <-- Inductive h. + 133*12^2k-1

OpenStudy (anonymous):

when you assume true for n=k. You should write a formula. It's much clearer than just writing an expression. For example: \[11^{k+1} +12^{2k-1}=133m where m is an integer.

OpenStudy (anonymous):

\[11^{k+1} +12^{2k-1}=133m\] where m is an integer.

OpenStudy (anonymous):

So when it comes to proving for n=k+!, you can write down LHS [Left Hand SIde]=

OpenStudy (anonymous):

and then go down the page until you reach the last line that you do which will be: \[=133[11m+{12^2k-1}]\] \[=133M\] where M [This M is different to the m before!!] is an integer.

OpenStudy (anonymous):

n=k+1*

OpenStudy (anonymous):

\[=133[11m+12^{2k-1}\]*

OpenStudy (anonymous):

Sorry about the typo there with the second last line. The website is a bit slow when it comes to typing.

OpenStudy (anonymous):

]**

OpenStudy (anonymous):

can you show me the first two line of the last step?

OpenStudy (anonymous):

Okay well you wrote what n=k was. You just substitute n for k. You would do the same thing for n=k+!. Your original equation was: \[11^{n+1} +12^{2n-1}=133m\] Then your very first step when proving n=k+! would be basically substituting n with k+1: \[LHS=11^{(k+1)+1}+12^{2(k+1)-1}\]

OpenStudy (anonymous):

n=k+1*

OpenStudy (anonymous):

then your second line would be trying to take out the extra terms so that you can put it in the form of n=k: which was this- \[LHS=11^{k+1} +12^{2k-1}\] SO your SECOND step would be to follow your first step: \[=(11\times 11^{k+1})+(12^2\times12^{2k-1})\]

OpenStudy (anonymous):

You're trying to put it in the form of n=k. That's the first two lines. Now you would be "rigging" the equation until you get it in the same form of 133m. And the last two lines I showed you.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No worries.